Consider the equation below:
2H2O → 2H2 + O2
a). How many moles of oxygen will be produced from 6.2 moles of water?
b). How many moles of water will be required to make 19.2 moles of oxygen?
2). The equation below shows how Ammonia gas is produced in the industry from Nitrogen and Hydrogen gas.
N2 + 3H2 <-------> 2NH3
a). How many moles of Ammonia can be produced from 5.6 moles of Nitrogen according to the equation?
b). How many moles of Hydrogen will be needed to react with 5.6moles of Nitrogen?

3). During a lab session, Iron fillings were made to react with excess hydrochloric acid to produce Iron (II) chloride and hydrogen gas.
Write a balanced equation for the reaction
If 5.6g of Iron fillings were used in the reaction, calculate:
The mass of Iron(II) chloride formed
The volume of hydrogen gas produced at s.t.p.
(Given: 1 mol of any gas occupies a volume of 22.4dm3 at s.t.p., Fe =56, Cl = 35.5, & H = 1)
1
Expert's answer
2018-05-14T08:14:24-0400
1a) n(H2O) = 6.2 mol

n(O2) = 6.2/2 = 3.1 mol

1b) n(O2) = 19.2 mol

n(H2O) = 19.2*2 = 38.4 mol

2a) n(N2) = 5.6 mol

n(NH3) = 2*5.6 = 11.2 mol

2b) n(N2) = 5.6 mol

n(H2) = 5.6*3 = 16.8 mol

3) Fe + 2HCl → FeCl2 + H2↑

m(Fe) = 5.6 g

n(Fe) = m/M = 5.6g/56g/mol = 0.1 mol

n(Fe) = n(FeCl2) = n(H2) = 0.1 mol

M(FeCl2) = 56 + 35.5*2 = 127 g/mol

m(FeCl2) = n*M = 0.1mol*127g/mol = 12.7 g

V(H2) = n*VM = 0.1mol*22.4dm3/mol = 2.24 dm3